How many liters of distilled water must be added to 140 liters of an 80 percent alcohol solution to obtain a 70 percent solution?
Answer by mananth(16946)

------ Percent ----------------quantity
Alcohol 80----------------140
water 0----------------x
Total 70----------------140+x
80*140 +0*x=70(140+x)
11200=9800+70x
-70x=9800-11200
-70x=-1400
/-70
x=20 gallons water